# revtex4-2Repair the float

### v2

@@ C.2 Third order terms with two Pauli matrices and one derivative
$\text{Tr}\partial_{k}Q^{\dagger}Q^{\dagger}\sigma_{j}Q^{\dagger}\sigma_{i}=\text{Tr}\tau_{3}\rho_{2}\partial_{k}Q^{\dagger}\tau_{3}\rho_{2}Q^{\dagger}\tau_{3}\rho_{2}\sigma_{j}\tau_{3}\rho_{2}Q^{\dagger}\tau_{3}\rho_{2}\sigma_{i}=-\text{Tr}\partial_{k}\tilde{Q}\tilde{Q}\sigma_{j}\tilde{Q}\sigma_{i}=-\text{Tr}\sigma_{j}\tilde{Q}\sigma_{i}\tilde{Q}\partial_{k}\tilde{Q},$ (40)
that is $\tilde{\gamma}_{lk}=-\tilde{\gamma}_{lk}^{*}$. We may thus define $\tilde{\gamma}_{lk}=i\gamma_{lk}$ where $\gamma_{lk}$ is a real second pseudotensor.
The contribution to the bulk Usadel equation is
$i\varepsilon_{ijl}\gamma_{lk}[\sigma_{j}g\partial_{k}g\sigma_{i}+\partial_{k}g\sigma_{i}g\sigma_{j}-\sigma_{j}\partial_{k}g\sigma_{i}g-\sigma_{i}g\sigma_{j}\partial_{k}g,g].$ (41)
In the normal state this becomes
$2\gamma_{lk}\tau_{3}\{\partial_{k}F,\sigma_{c}\}.$ (42)
@@
$J_{k}=i\varepsilon_{ijl}\gamma_{lk}[\sigma_{i}g\sigma_{j}g,g].$ (43)
In the normal state the Keldysh component of this expression becomes
$J_{k}^{K}=2i\varepsilon_{ijl}\tau_{3}\gamma_{lk}(\sigma_{i}F\sigma_{j}-\{\sigma_{i}\sigma_{j},F\}).$ (44)
Thus,
$\text{Tr}J_{k}^{K}=4\gamma_{lk}\text{Tr}\tau_{3}\sigma_{l}F,$ (45)
@@ C.3 Third order terms with one Pauli matrix and two derivatives
There exists two such terms; one with two $Q$’s:$\gamma_{ajk}\sigma_{i}\partial_{j}Q\partial_{k}Q$ and one with three $Q$’s: $\gamma_{ijk}\sigma_{i}Q\partial_{j}Q\partial_{k}Q$. Charge conjugation on the first term implies
$-\partial_{k}Q^{T}\partial_{j}Q^{T}\sigma_{y}\sigma_{i}\sigma_{y}=-\sigma_{y}\rho_{1}\tau_{1}\partial_{k}Q^{T}\tau_{1}\rho_{1}\rho_{1}\tau_{1}\partial_{j}Q^{T}\tau_{1}\rho_{1}\sigma_{y}\sigma_{i}=-\text{Tr}\partial_{k}Q\partial_{j}Q\sigma_{i}=-\text{Tr}\sigma_{i}\partial_{k}Q\partial_{j}Q,$ (46)
that is, $\tilde{\varkappa}_{ijk}=-\tilde{\varkappa}_{ikj}$ and we may write it as $i\frac{D}{4}\varepsilon_{jkl}\varkappa_{il}$, where $\varepsilon_{jkl}$ is the third rank antisymmetric pseudotensor and $\varkappa_{il}$ is a second rank tensor. The factor $\frac{D}{4}$ was chosen so that the $\chi_{al}$ corresponds to the spin-swapping coefficient in isotropic materials, as shown in section C.3.1. chronological symmetry gives
$\partial_{k}Q^{\dagger}\partial_{j}Q^{\dagger}\sigma_{i}=\tau_{3}\rho_{2}\partial_{k}Q^{\dagger}\rho_{2}\tau_{3}\tau_{3}\rho_{2}\partial_{j}Q^{\dagger}\rho_{2}\tau_{3}\sigma_{i}=\partial_{k}\tilde{Q}\partial_{j}\tilde{Q}\sigma_{i}=\sigma_{i}\partial_{k}\tilde{Q}\partial_{j}\tilde{Q},$ (47)
that is, $\tilde{\varkappa}_{ijk}=\tilde{\varkappa}_{ikj}^{*}=-\tilde{\varkappa}_{ijk}^{*}$. Hence $\varkappa_{il}$ is a real second rank tensor.
The contribution to the bulk equation is
$i\frac{D}{4}\varepsilon_{jkl}\varkappa_{il}[-\sigma_{i}\partial_{k}\partial_{j}g-\partial_{k}\partial_{j}g\sigma_{i},g]=0$ (48)
@@
$J_{k}=i\frac{D}{4}\varkappa_{il}(\varepsilon_{jki}[\sigma_{i}\partial_{j}g,g]+\varepsilon_{kji}[\partial_{j}g\sigma_{i},g])=i\frac{D}{4}\varkappa_{il}\varepsilon_{jki}[[\sigma_{i},\partial_{j}g],g].$ (49)
Note that for the second term we need to swap the $j,k$ indices to get a current contribution to $J_{k}$ and not $J_{j}$. In the normal state this expression becomes
$J_{k}^{K}=-iD\varkappa_{il}[\sigma_{i},\partial_{j}F]$ (50)
@@
$=-\text{Tr}\text{Tr}\partial_{k}Q\partial_{j}QQ\sigma_{i}=-\text{Tr}\sigma_{i}Q\partial_{k}Q\partial_{j}Q.$ (51)
Thus, the pseudotensor of this term satisfies $\tilde{\theta}_{ijk}=-\tilde{\theta}_{ikj}$ and we may write the tensor belonging to this term as $-\frac{D}{4}\varepsilon_{jkl}\theta_{il}$, where $\theta_{il}$ is a second rank tensor. The choice for the factor $-\frac{D}{4}$ was made to ensure that that in an isotropic material our tensor $\theta_{il}$ reduces to the conventional spin-Hall angle, as shown in section C.3.1.
@@
$\text{Tr}\partial_{k}Q^{\dagger}\partial_{j}Q^{\dagger}Q^{\dagger}\sigma_{i}=\text{Tr}\rho_{2}\tau_{3}\partial_{k}Q^{\dagger}\tau_{3}\rho_{2}\rho_{2}\tau_{3}\partial_{j}Q^{\dagger}\tau_{3}\rho_{2}\rho_{2}\tau_{3}Q^{\dagger}\tau_{3}\rho_{2}\sigma_{i}=-\text{Tr}\partial_{k}\tilde{Q}\partial_{j}\tilde{Q}\tilde{Q}\sigma_{i}=\text{Tr}\sigma_{i}\partial_{k}\tilde{Q}\partial_{j}\tilde{Q}\tilde{Q}=\text{Tr}\sigma_{i}\tilde{Q}\partial_{k}\tilde{Q}\partial_{j}\tilde{Q},$ (52)
and hence $\tilde{\theta}_{ijk}=-\tilde{\theta}_{ikj}^{*}=\tilde{\theta}_{ijk}^{*}$, it is real. Hence $\theta_{il}$ is a real second rank tensor. Its contribution to the bulk equation is
$-\frac{D}{4}\varepsilon_{jkl}\theta_{il}[\partial_{j}g\partial_{k}g\sigma_{i}-\partial_{k}g\sigma_{i}\partial_{j}g-\sigma_{i}\partial_{k}g\partial_{j}g,g],$ (53)
where it was used that the $\partial_{j}\partial_{k}g$-terms drop out due to antisymmetry in $j,k$. In the normal state this term drops out of the Usadel equation, since only one of the derivative terms can act on a Keldysh component while the retarded and advanced components are constant in the normal state. The contribution to the current is
$J_{k}=-\frac{D}{4}\theta_{il}(\varepsilon_{jkl}[\sigma_{i}g\partial_{j}g,g]+\varepsilon_{kjl}[\partial_{j}g\sigma_{i}g,g])=-\frac{D}{4}\varepsilon_{jkl}\theta_{il}[[\sigma_{i}g,\partial_{j}g],g].$ (54)
In the normal state, where $g^{K}=2\tau_{3}F$ this becomes
$J_{k}^{K}=\varepsilon_{jkl}D\theta_{il}\tau_{3}\{\sigma_{i},\partial_{j}F\}.$ (55)
@@ C.3.1 Isotropic materials
$J_{k}^{K}=-2D\text{Tr}\partial_{k}F+D\theta\varepsilon_{jki}\tau_{3}\{\sigma_{i},\partial_{j}F\}-iD\varkappa\varepsilon_{jki}[\sigma_{i},\partial_{j}F].$ (56)
@@
$\text{Tr}\sigma_{l}D\theta\varepsilon_{jki}\tau_{3}\{\sigma_{i},\partial_{j}F\}=2D\theta\varepsilon_{jki}\delta_{il}\text{Tr}\tau_{3}\sigma_{l}\partial_{j}f=-\theta\varepsilon_{jki}j_{ji}^{(0)},$ (57)
@@
$-iD\varkappa\varepsilon_{jki}\text{Tr}\sigma_{l}[\sigma_{i},\partial_{j}F]=-2iD\varkappa\varepsilon_{jki}i\varepsilon_{lim}\text{Tr}\sigma_{m}\partial_{j}F=-2D\varkappa\varepsilon_{jki}\varepsilon_{ilm}\text{Tr}\sigma_{m}\partial_{j}F=\varkappa\varepsilon_{jki}\varepsilon_{ilm}j_{jm}^{(0)},$ (58)
@@
$j_{kl}=\text{Tr}\sigma_{l}J_{k}^{K}=j_{kl}^{(0)}-\theta\varepsilon_{jki}j_{ji}^{(0)}+\varkappa(j_{lk}^{(0)}-\delta_{kl}j_{mm}^{(0)}),$ (59)
@@
Appendix D Vortex in a gyrotropic superconductor

### v3

@@ C.2 Third order terms with two Pauli matrices and one derivative
$\text{Tr}\partial_{k}Q^{\dagger}Q^{\dagger}\sigma_{j}Q^{\dagger}\sigma_{i}=\text{Tr}\tau_{3}\rho_{2}\partial_{k}Q^{\dagger}\tau_{3}\rho_{2}Q^{\dagger}\tau_{3}\rho_{2}\sigma_{j}\tau_{3}\rho_{2}Q^{\dagger}\tau_{3}\rho_{2}\sigma_{i}=-\text{Tr}\partial_{k}\tilde{Q}\tilde{Q}\sigma_{j}\tilde{Q}\sigma_{i}=-\text{Tr}\sigma_{j}\tilde{Q}\sigma_{i}\tilde{Q}\partial_{k}\tilde{Q}\;,$ (S49)
that is, $\gamma_{lk}$ is a real second pseudotensor.
The contribution to the bulk Usadel equation can be calculated using variation again
$-\frac{i}{8}\varepsilon_{ijl}\gamma_{lk}\Big{(}\sigma_{i}[\delta g,g]\sigma_{j}g\partial_{k}g+\sigma_{i}g\sigma_{j}[\delta g,g]\partial_{k}g+\sigma_{i}g\sigma_{j}g\partial_{k}[\delta g,g]\Big{)}$ (S50)
$=-\frac{i}{8}\varepsilon_{ijl}\gamma_{lk}\Big{(}g\sigma_{j}g\partial_{k}g\sigma_{i}-\sigma_{j}g\partial_{k}g\sigma_{i}g+g\partial_{k}g\sigma_{i}g\sigma_{j}-\partial_{k}g\sigma_{i}g\sigma_{j}g+\partial_{k}g\sigma_{i}g\sigma_{j}g-\sigma_{i}g\sigma_{j}g\partial_{k}g\Big{)}\delta g$
$-\frac{i}{8}\varepsilon_{ijl}\gamma_{lk}\Big{(}g\sigma_{i}g\sigma_{j}g-\sigma_{i}g\sigma_{j}\Big{)}\partial_{k}g\;.$ (S51)
In the first bracket, the fourth and fifth term cancel. The first and sixth term can be written as a commutator with $\sigma_{i}$, while the second and third can be written as a commutator with $\sigma_{j}$. After integration by parts, and using that each term changes sign upon interchanging $i,j$, we obtain
$-\frac{i}{8}\varepsilon_{ijl}\gamma_{lk}\Big{(}[g\partial_{k}g\sigma_{i}g-g\sigma_{i}g\partial_{k}g,\sigma_{j}]\Big{)}-\partial_{k}(\frac{i}{8}\varepsilon_{ijl}\gamma_{lk}[\sigma_{i}g\sigma_{j}g,g])\;.$ (S52)
@@
$J_{k}=\frac{i}{8}\varepsilon_{ijl}\gamma_{lk}[\sigma_{i}g\sigma_{j}g,g]=\frac{i}{16}\varepsilon_{ijl}\gamma_{lk}\{\sigma_{j}+g\sigma_{j}g,[\sigma_{i},g]\}\;.$ (S53)
and the torque of this term is
$\mathcal{T}=-\frac{i}{8}\varepsilon_{ijl}\gamma_{lk}[\{\partial_{k}g,g\sigma_{i}g\},\sigma_{j}]\;.$ (S54)
These two last equations correspond to the contribution proportional to the pseudotensor $\gamma_{ij}$ in the expressions Eqs. (8-9) of the main text.
@@ C.3 Third order terms with one Pauli matrix and two derivatives
There exists two such terms; one with two $Q$’s:
$-i\frac{D}{4}\varkappa_{ijk}\sigma_{i}\partial_{j}Q\partial_{k}Q\;,$ (S55)
and one with three $Q$’s:
$i\frac{D}{4}\theta_{ijk}\sigma_{i}Q\partial_{j}Q\partial_{k}Q\;.$ (S56)
The factors $\frac{D}{4}$ are chosen so that the $\chi_{al}$ corresponds to the spin-Hall and spin-swapping coefficients in the normal state, as shown in section C.3.1. Charge conjugation on the first term implies
$-\partial_{k}Q^{T}\partial_{j}Q^{T}\sigma_{y}\sigma_{i}\sigma_{y}=-\sigma_{y}\rho_{1}\tau_{1}\partial_{k}Q^{T}\tau_{1}\rho_{1}\rho_{1}\tau_{1}\partial_{j}Q^{T}\tau_{1}\rho_{1}\sigma_{y}\sigma_{i}=-\text{Tr}\partial_{k}Q\partial_{j}Q\sigma_{i}=-\text{Tr}\sigma_{i}\partial_{k}Q\partial_{j}Q\;,$ (S57)
that is, $\varkappa_{ijk}=-\varkappa_{ikj}$ and hence we may write it as $\varepsilon_{jkl}\varkappa_{il}$, where $\varepsilon_{jkl}$ is the third rank antisymmetric pseudotensor and $\varkappa_{il}$ is a second rank tensor. chronological symmetry gives
$\partial_{k}Q^{\dagger}\partial_{j}Q^{\dagger}\sigma_{i}=\tau_{3}\rho_{2}\partial_{k}Q^{\dagger}\rho_{2}\tau_{3}\tau_{3}\rho_{2}\partial_{j}Q^{\dagger}\rho_{2}\tau_{3}\sigma_{i}=\partial_{k}\tilde{Q}\partial_{j}\tilde{Q}\sigma_{i}=\sigma_{i}\partial_{k}\tilde{Q}\partial_{j}\tilde{Q}\;,$ (S58)
that is, $\varkappa_{il}$ needs to be a real second rank tensor.
The contribution to the bulk equation can be found using variation:
$=-i\frac{D}{4}\varepsilon_{jkl}\varkappa_{il}\text{Tr}\Bigg{(}\sigma_{i}\partial_{j}[\delta g,g]\partial_{k}g+\sigma_{i}\partial_{j}g\partial_{k}[\delta g,g]\Bigg{)}$ (S59)
$=i\frac{D}{4}\varepsilon_{jkl}\varkappa_{il}\text{Tr}\Bigg{(}\sigma_{i}\partial_{k}[\delta g,g]\partial_{j}g-\sigma_{i}\partial_{j}g\partial_{k}[\delta g,g]\Bigg{)}$ (S60)
$=i\frac{D}{4}\varepsilon_{jkl}\varkappa_{il}\text{Tr}\Bigg{(}\Big{(}\partial_{k}g\partial_{j}g\sigma_{i}-\partial_{j}g\sigma_{i}\partial_{k}g-\partial_{k}g\sigma_{i}\partial_{j}g+\sigma_{i}\partial_{j}g\partial_{k}g\Big{)}\delta g+\Big{(}g\partial_{j}g\sigma_{i}-\partial_{j}g\sigma_{i}g-g\sigma_{i}\partial_{j}g+\sigma_{i}\partial_{j}gg\Big{)}\partial_{k}\delta g\Bigg{)}$ (S61)
$=\text{Tr}\Bigg{(}\Big{(}i\frac{D}{4}\varepsilon_{jkl}\varkappa_{il}[\sigma_{i},\partial_{j}g\partial_{k}g]-\partial_{k}(i\frac{D}{4}\varepsilon_{jkl}\varkappa_{il}[[\sigma_{i},\partial_{j}g],g])\Big{)}\delta g\Bigg{)}\;.$ (S62)
For constant $\varkappa_{il}$ this reduces to
$0\;.$ (S63)
@@
$J_{k}=i\frac{D}{4}\varkappa_{il}(\varepsilon_{jki}[\sigma_{i}\partial_{j}g,g]+\varepsilon_{kji}[\partial_{j}g\sigma_{i},g])=i\frac{D}{4}\varkappa_{il}\varepsilon_{jkl}[[\sigma_{i},\partial_{j}g],g]\;.$ (S64)
Note that we need to swap the $j,k$ indices for the second term to get a current contribution to $J_{k}$ and not $J_{j}$. Because the contribution to the bulk equation vanishes the torque satisfies
$\mathcal{T}=i\frac{D}{4}\varepsilon_{jkl}\varkappa_{il}[\sigma_{i},\partial_{j}g\partial_{k}g]=i\frac{D}{4}\varkappa_{il}\epsilon_{jki}\partial_{k}[[\sigma_{i},\partial_{j}g],g]\;.$ (S65)
@@
$=-\text{Tr}\text{Tr}\partial_{k}Q\partial_{j}QQ\sigma_{i}=-\text{Tr}\sigma_{i}Q\partial_{k}Q\partial_{j}Q\;.$ (S66)
Thus, the pseudotensor of this term satisfies $\theta_{ijk}=-\theta_{ikj}$ and we may write the tensor belonging to this term as $\varepsilon_{jkl}\theta_{il}$, where $\theta_{il}$ is a second rank tensor. The choice for the factor $\frac{D}{4}$ was made to ensure that that in an isotropic material our tensor $\theta_{il}$ reduces to the conventional spin-Hall angle, as shown in section C.3.1.
@@
$\text{Tr}\partial_{k}Q^{\dagger}\partial_{j}Q^{\dagger}Q^{\dagger}\sigma_{i}=\text{Tr}\rho_{2}\tau_{3}\partial_{k}Q^{\dagger}\tau_{3}\rho_{2}\rho_{2}\tau_{3}\partial_{j}Q^{\dagger}\tau_{3}\rho_{2}\rho_{2}\tau_{3}Q^{\dagger}\tau_{3}\rho_{2}\sigma_{i}=-\text{Tr}\partial_{k}\tilde{Q}\partial_{j}\tilde{Q}\tilde{Q}\sigma_{i}=\text{Tr}\sigma_{i}\partial_{k}\tilde{Q}\partial_{j}\tilde{Q}\tilde{Q}=\text{Tr}\sigma_{i}\tilde{Q}\partial_{k}\tilde{Q}\partial_{j}\tilde{Q}\;,$ (S67)
and hence $\theta_{il}$ is a real second rank tensor. Its contribution to the bulk equation can be found using
$\frac{D}{4}\varepsilon_{jkl}\theta_{il}\sigma_{i}\Big{(}[\delta g,g]\partial_{j}g\partial_{k}g+g\partial_{j}[\delta g,g]\partial_{k}g+g\partial_{j}g\partial_{k}[\delta g,g]\Big{)}$ (S68)
$\frac{D}{4}\varepsilon_{jkl}\theta_{il}\sigma_{i}\Big{(}[\delta g,g]\partial_{j}g\partial_{k}g-g\partial_{k}[\delta g,g]\partial_{j}g+g\partial_{j}g\partial_{k}[\delta g,g]\Big{)}$ (S69)
$=\frac{D}{4}\varepsilon_{jkl}\theta_{il}\Bigg{(}\Big{(}g\partial_{j}g\partial_{k}g\sigma_{i}-\partial_{j}g\partial_{k}g\sigma_{i}g-\partial_{k}g\partial_{j}g\sigma_{i}g+\partial_{j}g\sigma_{i}g\partial_{k}g+\partial_{k}g\sigma_{i}g\partial_{j}g-\sigma_{i}g\partial_{j}g\partial_{k}g\Big{)}\delta g$
$+\Big{(}-g\partial_{j}g\sigma_{i}g+\partial_{j}g\sigma_{i}+g\sigma_{i}g\partial_{j}g+\sigma_{i}\partial_{j}g\Big{)}\partial_{k}g\Bigg{)}$ (S70)
$=-\Bigg{(}\frac{D}{4}\varepsilon_{jkl}\theta_{il}\Big{(}[\sigma_{i},g\partial_{j}g\partial_{k}g]\Big{)}+\partial_{k}\Big{(}\frac{D}{4}\varepsilon_{jkl}\theta_{il}\{\sigma_{i}+g\sigma_{i}g,\partial_{j}g\}\Big{)}\Bigg{)}\;,$ (S71)
where it was used that in the first bracket the second and the third term cancel, just like the fourth and the fifth, due to the antisymmetry in $j,k$.
In the normal state this term drops out of the Usadel equation, since only one of the derivative terms can act on a Keldysh component while the retarded and advanced components are constant in the normal state. The contribution to the current is
$J_{k}=-\frac{D}{4}\theta_{il}(\varepsilon_{jkl}[\sigma_{i}g\partial_{j}g,g]+\varepsilon_{kjl}[\partial_{j}g\sigma_{i}g,g])=-\frac{D}{4}\varepsilon_{jkl}\theta_{il}[[\sigma_{i}g,\partial_{j}g],g]=\frac{D}{4}\varepsilon_{jkl}\theta_{il}\{\sigma_{i}+g\sigma_{i}g,\partial_{j}g\}\;.$ (S72)
Thus, the torque is given by
$\mathcal{T}=-\frac{D}{4}\epsilon_{jkl}\theta_{il}[\sigma_{i},g\partial_{j}g\partial_{k}g]\;.$ (S73)
@@ C.3.1 Isotropic materials
$J_{k}^{K}=-2D\text{Tr}\partial_{k}F+D\theta\varepsilon_{jki}\tau_{3}\{\sigma_{i},\partial_{j}F\}-iD\varkappa\varepsilon_{jki}[\sigma_{i},\partial_{j}F]\;.$ (S74)
@@
$\text{Tr}\sigma_{l}D\theta\varepsilon_{jki}\tau_{3}\{\sigma_{i},\partial_{j}F\}=2D\theta\varepsilon_{jki}\delta_{il}\text{Tr}\tau_{3}\sigma_{l}\partial_{j}f=-\theta\varepsilon_{jki}j_{ji}^{(0)}\;,$ (S75)
@@
$-iD\varkappa\varepsilon_{jki}\text{Tr}\sigma_{l}[\sigma_{i},\partial_{j}F]=-2iD\varkappa\varepsilon_{jki}i\varepsilon_{lim}\text{Tr}\sigma_{m}\partial_{j}F=-2D\varkappa\varepsilon_{jki}\varepsilon_{ilm}\text{Tr}\sigma_{m}\partial_{j}F=\varkappa\varepsilon_{jki}\varepsilon_{ilm}j_{jm}^{(0)}\;,$ (S76)
@@
$j_{kl}=\text{Tr}\sigma_{l}J_{k}^{K}=j_{kl}^{(0)}-\theta\varepsilon_{jki}j_{ji}^{(0)}+\varkappa(j_{lk}^{(0)}-\delta_{kl}j_{mm}^{(0)})\;,$ (S77)
@@
Appendix D Keldysh structure and kinetic equation
In this section we give a brief outline of the Keldysh-structure of the Green’s functions used in this letter, based on [33], and present the kinetic equations in the normal state.
[…]
It is convenient to write these two different branches in a matrix structure, that is, to define the Green’s function on the contour as
$\tilde{g}(t,t^{\prime})=\begin{bmatrix}\tilde{g}(t^{+},t^{\prime+})&\tilde{g}(t^{+},t^{\prime-})\\ \tilde{g}(t^{-},t^{\prime+})&\tilde{g}(t^{-},t^{\prime-})\end{bmatrix}\;,$ (S78)
where $\pm$ refers to the positive and negative branch respectively. For causal theories, the entries of Eq. (S78) are not independent. Indeed, if $t>t^{\prime}$, one may traverse the track from $t$ to $\infty$ and back. Since this concerns only future times, by causality, no physical observable can change and hence $\tilde{g}(t^{+},t^{\prime\pm})=\tilde{g}(t^{-},t^{\prime\pm})$. By an entirely similar argument, if $t<t^{\prime}$, we find $\tilde{g}(t^{\pm},t^{\prime+})=\tilde{g}(t^{\pm},t^{\prime-})$.
To exploit this dependence, it is convenient to rotate the Keldysh space by
$g=L\rho_{3}\tilde{g}L^{\dagger},$ (S79)
where $L=1/\sqrt{2}(1-i\rho_{2})$ and $\rho_{i},i=1,2,3$ denote the Pauli matrices in Keldysh space. In this space, the Green’s function takes the form
$g=\begin{bmatrix}g^{R}&g^{K}\\ 0&g^{A}\end{bmatrix}\;,$ (S80)
where $g^{R,A,K}$ are the retarded, advanced and Keldysh Green’s functions respectively. The retarded and advanced Green’s functions describe equilibrium properties, such as the density of states. The Keldysh Green’s function may be used to describe nonequilibrium physics and may be parametrized as $g^{K}=g^{R}F-Fg^{A}$, where $F$ represents the distribution function of the system. Here $F$ is diagonal in Nambu space, but may have any structure in spin space.
The Green’s function satisfies the Usadel equation, the saddle point equation of the NLSM. In the absence of spin-Hall and spin-swapping terms these correspond to Eqs. 7-9 in the main text. In the presence of spin-Hall and spin-swapping terms, current and torque are modified to
$\mathcal{J}_{k}$ $=-Dg\partial_{k}g+i\alpha_{kj}g[\sigma_{k},g]+\frac{i}{16}\epsilon_{ijl}\gamma_{lk}\{[\sigma_{i},g],\sigma_{j}+g\sigma_{j}g\}+i\frac{D}{4}\varkappa_{il}\epsilon_{jkl}[[\sigma_{i},\partial_{j}g],g]+\frac{D}{4}\epsilon_{jkl}\theta_{il}\{\sigma_{i}+g\sigma_{i}g,\partial_{j}g\},$ (S81)
$\mathcal{T}$ $=-i\alpha_{kj}[\sigma_{k},g\partial_{j}g]-\frac{i}{8}\varepsilon_{ijl}\gamma_{lk}[\{\partial_{k}g,g\sigma_{i}g\},\sigma_{j}]-i\frac{D}{4}\varkappa_{il}\epsilon_{jkl}\partial_{k}([[\sigma_{i},\partial_{j}g],g])-\frac{D}{4}\epsilon_{jkl}\theta_{il}[\sigma_{i},g\partial_{j}g\partial_{k}g]\,.$ (S82)
In this case, pair amplitudes are absent while the density of states is constant within the quasiclassical formalism. Thus, the retarded and advanced Green’s function are equal to $\tau_{3}\delta(t-t^{\prime})$ and $-\tau_{3}\delta(t-t^{\prime})$ respectively, independent of spatial coordinates.
However, the Keldysh equation remains nontrivial, because the distribution functions may depend on space, as happens when we apply a voltage. By substitution of $g^{R}=-g^{A}=\tau_{3}\delta(t-t^{\prime})$ in these equations, we obtain a linear equation for $g^{K}$:
$\partial_{k}J_{k}^{K}$ $=-\tau_{3}\partial_{t}g^{K}-i\alpha_{kj}\tau_{3}[\sigma_{j},\partial_{k}g^{K}]-\frac{i}{4}\gamma_{lk}(\varepsilon_{ijl}\sigma_{i}\partial_{k}g^{K}\sigma_{j}+i\{\sigma_{l},g^{K}\}))\;,$ (S83)
$J^{K}_{k}$ $=-D\tau_{3}\partial_{k}g^{K}+i\alpha_{kj}\tau_{3}[\sigma_{j},g^{K}]+\frac{i}{4}\gamma_{lk}(\varepsilon_{ijl}\sigma_{i}g^{K}\sigma_{j}-i\{\sigma_{l},g^{K}\})-i\frac{D}{2}\varkappa_{il}\epsilon_{jkl}\tau_{3}[\sigma_{i},\partial_{j}g^{K}]+\frac{D}{2}\epsilon_{jkl}\theta_{il}\{\sigma_{i},\partial_{j}g^{K}\}\;.$ (S84)
These equations may also be expressed in terms of the distribution functions, which in the normal state are related to the Keldysh component of the Green’s function via $g^{K}=2\tau_{3}F$, as:
$\partial_{k}J_{k}^{K}$ $=-2\partial_{t}F-2i\alpha_{kj}[\sigma_{j},\partial_{k}F]-\frac{i}{2}\gamma_{lk}\tau_{3}(\varepsilon_{ijl}\sigma_{i}\partial_{k}F\sigma_{j}+i\{\sigma_{l},F\}))\;,$ (S85)
$J^{K}_{k}$ $=-2D\partial_{k}F+i\alpha_{kj}[\sigma_{j},F]+\frac{i}{2}\gamma_{lk}\tau_{3}(\varepsilon_{ijl}\sigma_{i}F\sigma_{j}-i\{\sigma_{l},F\})-iD\varkappa_{il}\epsilon_{jkl}[\sigma_{i},\partial_{j}F]+D\epsilon_{jkl}\theta_{il}\tau_{3}\{\sigma_{i},\partial_{j}F\}\;.$ (S86)
[…]
$\partial_{k}j_{k}$ $=-\partial_{t}\mu\;,$ (S87)
$\partial_{k}j_{kl}$ $=-\partial_{t}S^{l}+\varepsilon_{lmr}h_{m}S^{r}+\gamma_{lm}\partial_{m}\mu-\frac{1}{2}\Gamma_{0}S^{l}+\frac{1}{2}\Gamma_{lm}S^{m}+\alpha_{km}\varepsilon_{lmr}\partial_{k}S^{r}\;,$ (S88)
$j_{k}$ $=-D(\partial_{k}\mu-\theta_{il}\epsilon_{jkl}\partial_{j}S^{i})+\gamma_{ik}S^{i}\;,$ (S89)
$j_{kl}$ $=-D(\partial_{k}S^{l}-\theta_{il}\epsilon_{jkl}\partial_{j}\mu-\varkappa_{ir}\epsilon_{jkr}\epsilon_{iml}\partial_{j}S^{m})-\alpha_{km}\varepsilon_{lmr}S^{r}\;,$ (S90)
where $\Gamma_{0}=\text{Tr}\Gamma_{ab}$. These equations can be supplemented by suitable boundary conditions that depend on the type of boundary chosen, for example for a boundary with vacuum $n_{k}J_{k}=0$ and $n_{k}J_{kl}=0$ for all $l$.
The first two equations are the diffusion equations for charge and spin, respectively. As expected, charge conservation leads to a zero torque in Eq. (S87). For the spin current, there are several types of torques, see Eq. (S88). First of all, even in the absence of spin-charge coupling, there may be the torque of an effective exchange field. $\gamma_{lm}\partial_{m}\mu$ is a generalization of the spin generation torque [65], it creates spin accumulation from electrical currents. On the other hand $-\frac{1}{2}\Gamma_{0}S^{l}+\frac{1}{2}\Gamma_{lm}S^{m}$ is a general spin relaxation term. For isotropic systems, $\Gamma_{lm}=\frac{1}{3}\Gamma_{0}\delta_{lm}$, and the torque reduces to $-\frac{1}{3}\Gamma_{0}S^{l}$. For anisotropic systems, spin-relaxation along one spin-axis may be different from that for other spin axes. Lastly, there is a torque due to spin precession. The spin-swapping and spin Hall terms do not give any torques in the normal state, even though their contribution to the torque is nonzero in the nonlinear Usadel equation.
Next to this, both the electrical current and spin current definitions are altered. The electrical current, shown in Eq. (S89) consists of three terms. First of all there is the usual diffusion term $-D\partial_{k}\mu$. Then there is the spin-Hall contribution $\theta_{jki}\partial_{i}S^{j}$, which is responsible for the inverse spin-Hall effect. Because $\theta_{jki}=\varepsilon_{jkl}\theta_{il}$ is necessarily antisymmetric in $i,k$, the generated electrical current is always perpendicular to the gradient of the spin. For isotropic materials $\theta_{jki}=\theta\epsilon_{jki}$ and hence the spin of the generated spin current is perpendicular to both the current direction and spin gradient. However, for certain crystal structures the spin may as well be parallel to either of these directions as well. Lastly, there is the term $\gamma_{ik}S^{i}$, gives a finite contribution to the electrical current in the presence of an excess spin and is therefore responsible for spin-galvanic effects.
The spin current, Eq. (S90), has four contributions. Next to the usual diffusion term $-D\partial_{k}S^{l}$, there is the spin-Hall term $-\theta_{lkj}\partial_{j}\mu$, which generates a spin current from a gradient in the chemical potential, i.e. it is responsible for the spin-Hall effect. By structure of the spin-Hall tensor, the spatial direction of this spin current is perpendicular to the gradient. The allowed directions of spin depends on the symmetries of the material. For isotropic materials, this spin is required to be perpendicular to both the gradient and spin current directions. Lastly, there exists a generalization of the spin swapping term [42] $\varkappa_{ijk}\varepsilon_{ilm}\partial_{j}S^{m}$, which mixes spin currents that have both perpendicular spin and perpendicular direction. The allowed form of the tensor, and hence the allowed relative orientation between direction of the spin and spatial direction of the spin current depends on the crystal symmetry. Specifically, for isotropic materials $\varkappa_{ijk}=\varkappa\epsilon_{ijk}$. Lastly, the spin-precession also alters the spin current.
The obtained spin diffusion equations contain all symmetry allowed terms, and are therefore completely general for normal metals with spin-charge coupling. The symmetry arguments only restrict the forms of the tensors, but do not relate the tensors of different terms. For this reason, the magnitudes of spin-precession, spin-relaxation and spin-generation are unrelated. For specific models, such as the SU(2) Rashba-like model [16] the different tensors are related. However, one should take into account that extrinsic spin-orbit coupling is always allowed and alters these relations.
Appendix E Vortex in a gyrotropic superconductor